OpenStudy (anonymous):

c=5/9(F-32) solve for F

OpenStudy (jamesj):

Multiply first both sides by 9/5 and you have \[{9 \over 5} C = {9 \over 5}{5 \over 9}(F - 32) = F - 32\] Hence F = ...?